OpenStudy (jordanloveangel):

Solve the equation for y. 8x – 9y = 11

OpenStudy (jordanloveangel):

i never understand how 2 do that

OpenStudy (anonymous):

D. Move 8x to the RHS and then divide each side by -9 .

OpenStudy (anonymous):

You have 8x -9y = 11. To preserve the equality : - when you add something to the left of the "=", you have to add it to the right - when you add something to the right of the "=", you have to add it to the left - when you subtract something from the left of the "=", you have to subtract it from the right - when you subtract something from the right of the "=", you have to subtract it from the left Therefore: - when you move something from the left side of the "=" to the right side of the "=", you always "change" signs. So if you have 2 - x = 5, you take "-x" and flip it to the right side with a changed sign. The equivalent of 2 - x =5 is thus 2 = 5 + x. You can view this as a simple "just adding x to both left and right sides" so whatever interpretation of this logical "phenomenon" you prefer, take it as that.

OpenStudy (jordanloveangel):

so d @AngusV

OpenStudy (anonymous):

As for this exercise - don't be scared of the two variables (x and y). If they (the new world government) are interested in the y, then treat x like the inferior and useless variable that it is. 8x - 9y = 11. Either you see it as fliping the worthless 8x or sending it to an extermination camp (because it's useless and we don't need it) either way the result will be: -9y = 11 - 8x Now to reveal the final solution for our great nation, we need to level the masses and give power to the fundamental class. y = (11 - 8x)/ (-9) Of course, we'll want everything to be in top shape so our propaganda needs to be perfect. We will draw the masses to focus on the 8x first and foremost - and with the help of propaganda, we have: y= 8x/9 - 11/9 That concludes my short presentation on fascism and equations.

OpenStudy (deadpoolroxz):

hmmm

OpenStudy (deadpoolroxz):

yeah hez right

OpenStudy (deadpoolroxz):

hey jessie

OpenStudy (jordanloveangel):

thx guys